Given this information:
Expected demand during lead time = 285 units
Standard deviation of lead time demand = 16 units
Use Table.
Determine each of the following, assuming that lead time demand is distributed normally:

a. The ROP that will provide a risk of stockout of 1 percent during lead time. (Round your answer to the nearest whole number.)

ROP _____ units
  
b. The safety stock needed to attain a 1 percent risk of stockout during lead time. (Do not round intermediate calculations. Round your answer to the nearest whole number.)

Safety stock ____ units

c-1.
Would a stockout risk of 4 percent require more or less safety stock than a 1 percent risk?

  • More

  • Less



c-2.
Would the ROP be larger, smaller, or unaffected if the acceptable risk were 4 percent instead of 1 percent?
  

  • Unaffected

  • Smaller

  • Larger
Answer #1

A. Service level = 100 - 1% = 99%

Z = 2.58 ( for 99% service level)

ROP = demand during lead time + ( Z * standard deviation of lead time)

= 285 + (2.58 * 16) = 326.28

B. Safety stock = Z * standard deviation of lead time

= 2.58 * 16 = 41.28

C. For stockout risk of 4 percent, Z value will be lower thus safety stock will be lower. The correct option is LESS.

D. Similarly ROP will be lower, the correct option is SMALLER.